The diagonal brace on a gate is 5 feet long. THe height of the gate is 4 feet long. How wide is the gate?
Answer by checkley77(12844)

x^2+4^2=5^2
x^2+16=25
x^2=25-19
x^2=9
x=sqrt9
x=3 ft is the width of the gate.